# Covariate Adjusted Response Adaptive Design with Delayed Outcomes

### v1

@@ 5 Synthetic case study
To evaluate the performance of the proposed forward-looking CARA design, we conduct a synthetic case study based on the randomized experiment reported in Fahey et al. (2020). Through this synthetic case study, we first demonstrate that both of the two forward-looking CARA designs (Section 3) deliver valid statistical inference (Table 1), supporting the theoretical result established in Theorem 2. Second, numerical evidence suggests that our fully forward-looking CARA design with the power maximization objective achieves a higher estimation efficiency compared to both the Neyman allocation and the complete randomization design in the presence of delayed responses (Figures 5 and 6), further validating the theoretical results in Theorem 5. Finally, compared to both the ethical design (Rosenberger et al., 2001) and complete randomization, our fully forward-looking CARA design can provide greater failure reduction, demonstrating its effectiveness in improving participant welfare while maintaining statistical power (Figure 5).
@@
The original experiment conducted by Fahey et al. (2020) aims to evaluate the effectiveness of cash transfers on retention in care among people living with HIV (PLHIV), using viral load as the primary response variable. The experiment participants were adults aged 18 years or older diagnosed with HIV, who were randomized to receive cash transfers (treatment arm) or no cash transfers (control arm). The response variable, viral load, was assessed six months after enrollment. Although the original study was not conducted in an adaptive fashion, adaptation can be applied since the participants were enrolled sequentially. However, a challenge is that the primary responses of the participants were collected with delays, with some responses collected longer than six months after the treatment assignment. As illustrated in Example 1, the empirically observed delay mechanism exhibits significant heterogeneity across arms and covariate strata. Our fully forward-looking CARA design can be utilized in such scenarios to help the experiment more effectively achieve its design goals. In this synthetic case study, we conduct a simulation experiment in which the underlying data-generating process — including the response delay mechanism, covariate distribution, and the (conditional) potential outcomes distributions — is calibrated with estimates from the (Fahey et al., 2020) study. We set $n_{t}=100$ for $t=1,\ldots,T$, and $T$ ranges in $\{4,6,8\}$.
The primary response variable is the viral load measured at the end of each stage. We consider two different transformations of the original viral load measurements, which shall be illustrated in the two simulation setups below. Delays in collecting the response variable are observed, and as discussed in Example 1, the delay mechanism depends on the arm and the biological sex of the participant. We consider two treatment arms: the control arm $A_{i}=0$, defined as “received no cash transfer”, and the treatment arm, $A_{i}=1$, defined as “received cash transfer.” Biological sex is included as the covariate $X_{i}$, where $X_{i}=1$ and 00 represent “male” and “female,” respectively. From the original experimental data, the distribution of biological sex consists of 36% male and 64% female, which we adopt as the data-generating process of $X_{i}$ in our Monte Carlo exercise. Specifically, we generate the covariate $X_{i}$ and treatment $A_{i}$ as follows:
$X_{i}$ $\sim\text{Bernoulli}(0.36),\ i=1,\ldots,N,$
$A_{i}|X_{i}=x,S_{i}=t$ $\sim\text{Bernoulli}(e_{t}(1|x)),\ t=1,\ldots,T,$
where $e_{t}(1|x)$ entails different designs we aim to compare. In what follows, we shall illustrate the design strategies in comparison under two sets of simulation setups tailored to each design objective:
Setup 1: Power maximization. For the power maximization design objective in Eq (2), we compare four designs: (1) the fully forward-looking CARA design for power maximization (Section 3); (2) the fully forward-looking CARA design further enhanced by the doubly adaptive biased coin (DBCD) method (Hu and Zhang, 2004; Tymofyeyev et al., 2007); (3) complete randomization; (4) treatment allocation based on Neyman allocation. Furthermore, in HIV research, it is standard practice to apply a log transformation to raw viral load measurements, given their wide range. Consequently, the response variable in this simulation setup is defined as: $\log(\texttt{viral load})$. We generate the response variable as
$Y_{i}|A_{i}=a,X_{i}=x\sim\mathcal{N}\left(\mu(x,a),\ \sigma^{2}(x,a)\right)$
to mimic the distribution of $\log(\texttt{viral load})$ in the original data, where for women $\mu(0,1)=2.50$, $\mu(0,0)=2.98$, $\sigma(0,1)=0.36$, $\sigma(0,0)=2.06$, and for men $\mu(1,1)=2.47$, $\mu(1,0)=2.72$, $\sigma(1,1)=0.82$, and $\sigma(1,0)=0.31$. Lastly, to evaluate the power of different designs under the power maximization objective, we vary the magnitude of ATE in our data-generating process by perturbing $\mu(1,0)$, such that the ATE $\tau(1,0)=0.33+\delta$, where $\delta\in\{-0,33,\ldots,0,\ldots,0.33\}$.
@@
Setup 2: Failure reduction. For the failure reduction design objective in Eq (3), we compare four designs: (1) the proposed fully forward-looking CARA design for failure reduction; (2) the proposed design enhanced by the DBCD method (Hu and Zhang, 2004); (3) the ethical design introduced by Rosenberger et al. (2001); (4) complete randomization. In this simulation setup, since the goal is to reduce the percentage of failures, we transform the original viral load measurements into a binary response variable, indicating whether the participant has achieved viral suppression. Specifically, if a participant’s viral load is less than $1,000$ copies per mL, we set $Y_{i}=1$, representing viral suppression. Otherwise, $Y_{i}$ is set to 00 (no viral suppression). We generate the binary response variable as
$Y_{i}|A_{i}=a,X_{i}=x\sim\text{Bernoulli}(\mu(x,a)),\ i=1,\ldots,N,$
where $\mu(0,1)=0.78$ and $\mu(0,0)=0.57$ represent the expected probabilities of viral suppression for females in the treatment and control arms, respectively. Similarly, $\mu(1,1)=0.84$ and $\mu(1,0)=0.63$ represent the expected probabilities of viral suppression probabilities for males in the treatment and control arms, respectively. Lastly, to evaluate the power of different designs under the failure reduction objective, we vary the magnitude of ATE in our data-generating process. For Setup 2, we vary the ATE value by perturbing $\mu(1,0)$, such that the ATE $\tau(1,0)=0.21+\delta$, where $\delta\in\{-0,21,\ldots,0,\ldots,0.21\}$.
In our synthetic case study, we incorporate both arm- and covariate-specific outcome delays following empirical estimates obtained in the original experimental data. We generate the delay variable $D_{i}$ from multinomial distribution:
$D_{i}|A_{i}=1,X_{i}=1$ $\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|1,1)\right),\ D_{i}|A_{i}=1,X_{i}=0\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|1,0)\right),$
$D_{i}|A_{i}=0,X_{i}=1$ $\sim\text{Multinomial}\left(1,\bm{\rho}(\cdot|0,1)\right),\ D_{i}|A_{i}=0,X_{i}=0\sim\text{Multinomial}(1,\bm{\rho}(\cdot|0,0)),$
where the first row represents the response delay mechanism for males and females under the treatment arm, while the second row represents the delay mechanism for males and females under the control arm. We slightly abuse the notation to define $\bm{\rho}(\cdot|a,x):=(\mathbb{P}(D_{i}=0|a,x),\ldots,\mathbb{P}(D_{i}=T-1|a,x))^{\intercal}$, where each element represents the probability that the response of subject $i$ is delayed by exactly $d$ stages, for $d=0,\ldots,T-1$. When $T=4$, the parameter values for the delay mechanisms are as follows: $\bm{\rho}(\cdot|1,1)=(0.64,0.18,0.07,0.03)^{\intercal}$, $\bm{\rho}(\cdot|1,0)=(0.63,0.18,0.05,0.02)^{\intercal}$, $\bm{\rho}(\cdot|0,1)=(0.55,0.23,0.10,0.02)^{\intercal}$, and $\bm{\rho}(\cdot|0,0)=(0.54,0.11,0.21,0.01)^{\intercal}$. We defer the details of the delay mechanism generation for $T=6$ and $T=8$ to the Supplementary Materials. To generate delayed responses, at each Stage $t$, we first generate responses for all $n_{t}$ subjects. Based on the delay mechanism, we then randomly select a subset of subjects with $(A_{i}=a,X_{i}=x)$ and label their responses as NA, indicating that their responses will be delayed to be observed.
For both simulation setups, we generate 500 Monte Carlo samples. To initialize the experiment, we set the treatment allocation to be $1/2$ in each covariate stratum in Stage 1 for all the designs. For Setup 1, we shall compare the variance of the ATE estimator under different design strategies in Figure 5 (A) – (C). For Setup 2, we shall compare the failure reduction rate under different design strategies in Figure 5 (D) – (E). For Setup 1 and Setup 2, we consider the conservative experimenter’s expectations about future delays. We defer the simulation results for neutral and optimistic experimenter’s expectations to Supplementary Materials.
For both setups, we also present the coverage probabilities (Table 1) and the power curve (Figure 6) to evaluate the statistical validity.
Figure 5: The variance of the ATE estimator under the power maximization design objective (A) – (C) and the failure reduction rate under the failure reduction design objective (D) – (F) across different design strategies.
Figure 6: The power curves under the (A) power maximization design objective and the (B) failure reduction design objective across various design strategies.
@@
Conservative 0.95 (0.01) 0.94 (0.01)
Optimistic 0.94 (0.01) 0.96 (0.01)
Neutral 0.95 (0.01) 0.96 (0.01)
@@
Conservative 0.96 (0.01) 0.95 (0.01)
Optimistic 0.95 (0.01) 0.94 (0.01)
Neutral 0.94 (0.01) 0.95 (0.01)
Table 1: Coverage probability comparison across different delay perspectives under two different design objectives.
Figure 5 (A) – (C) shows the variance of the ATE estimator under the power maximization objective across experimental stages for different designs. The results show that the proposed fully forward-looking CARA design and its DBCD-enhanced counterpart exhibit a smaller deviation from the oracle variance (indicated by the dashed line) than the complete randomization design and the Neyman allocation. This suggests that the proposed forward-looking design achieves higher estimation efficiency. Furthermore, Figure 5 (A) – (C) demonstrates that the efficiency of the fully forward-looking CARA design progressively improves as the experimental stages advance. We conjecture that this is because, as we obtain more information about the delay mechanism, the treatment allocation approaches the optimal allocation, leading to more significant changes over time. In contrast, while the Neyman allocation design has a notable improvement in efficiency after Stage 1, its variance stabilizes in subsequent stages. This is because, after the first stage, the Neyman allocation design adjusts for the $1/2$ treatment allocation initialized in Stage 1. Since Neyman allocation does not account for the delay mechanism, its variance quickly stabilizes over time.
Figure 5 (D) – (E) shows a similar trend under the failure reduction design objective. The fully forward-looking CARA design and its DBCD-enhanced counterpart achieve high failure reduction rates. Moreover, while the fully forward-looking CARA design shows a gradual decline in failure reduction as experimental stages increase, the failure reduction rates for the ethical design and the complete randomization design remain stable over time. This suggests that the adaptive nature of the fully forward-looking CARA design facilitates dynamically improved performance across stages. Lastly, our proposed fully forward-looking CARA design enhanced by DBCD overall exhibits a more robust performance. We conjecture that this is because the DBCD design targets the current stage treatment allocation toward the optimal treatment allocation, thereby further improving the empirical performance (i.e., estimation efficiency and failure rate reduction).
As our design is rigorously guided by the semiparametric efficiency theory to improve the estimation efficiency of the average causal effect, we present power curves for correctly detecting the true ATEs in Figure 6. Figure 6 suggests that when the ATE is 0, all considered designs control the type I error rate at the nominal level. Across two design objectives, the proposed forward-looking design overall demonstrates higher power in detecting the treatment effect compared to other designs, further validating the estimation efficiency of our proposed design.
To demonstrate the statistical validity of our proposed designs under different experimenters’ anticipation of delay mechanism, we present the coverage probabilities of the constructed confidence interval correctly covering the true ATE under both design objectives in Table 1. Table 1 shows that both the proposed fully forward-looking CARA design and its DBCD-enhanced design attain the nominal-level coverage probabilities across various delay perspectives. This suggests that, regardless of the perspective on delay when carrying out the experiment, the resulting ATE estimators under the proposed fully forward-looking design provide valid statistical inferences.
In summary, this synthetic case study demonstrates that our proposed fully forward-looking designs provide valid statistical inference in the presence of delayed responses. Additionally, our proposed designs exhibit higher estimation efficiency under the power maximization objective and improved failure reduction rate under the failure reduction objective compared to the competing designs, further validating the theoretical results presented in Section 4.
6 Discussion and concluding remarks
In this manuscript, we have proposed a novel fully forward-looking CARA design to address the challenges posed by delayed outcomes. By accounting for response delay mechanisms to identify optimal treatment allocations, our design achieves substantial improvements in experimental efficiency and participant welfare compared to traditional CARA designs. Through rigorous theoretical development, we derived semiparametric efficiency bounds that account for covariate-dependent delays and demonstrated that the proposed design achieves consistent and asymptotically normal treatment effect estimators. Our simulation studies further validated the practical advantages of this CARA design, showcasing its ability to outperform classical methods like Neyman allocation and complete randomization in terms of statistical efficiency and ethical outcomes. Notably, the forward-looking design dynamically adapts to evolving data, ensuring robust performance across varying experimental stages. Moreover, our work provides practical tools for implementing delay-informed CARA designs, offering pathways for future studies in clinical trials and other adaptive experimental settings.

### v2

@@ 5 Synthetic case study
To evaluate the performance of the proposed forward-looking CARA design, we conduct simulation studies using a data generation mechanism calibrated from the study of Fahey, Njau, Katabaro, Mfaume, Ulenga, Mwenda et al. (2020). Their original experiment aims to evaluate the effectiveness of cash transfers on retention in care among people living with HIV (PLHIV), using viral load as the primary response variable. The experiment participants were adults aged 18 years or older diagnosed with HIV, who were randomized to receive cash transfers (treatment arm, $A_{i}=1$) or no cash transfer (control arm, $A_{i}=0$). The original study involved two treatment arms with different levels of cash incentives, which are combined into one group in our synthetic case study. The response variable, viral load, was assessed six months after enrollment. Although the original study was not conducted in an adaptive fashion, adaptation can be applied since the participants were enrolled sequentially. However, a challenge is that the primary responses of the participants were collected with delays, with some responses collected longer than six months after the treatment assignment; see Figure 1. Our synthetic case study conducts a simulation experiment in which the underlying data-generating process (delay mechanism, covariate distribution, and potential outcome distributions) is calibrated with estimates from the original study. For the first set of simulation results, we consider biological sex as the only covariate $X_{i}$, leading to two subgroups: $X_{i}=\mathtt{M}$ and $X_{i}=\mathtt{F}$ for male and female participants. From the original data, the distribution of biological sex consists of 36% male and 64% female, which we adopt as the data-generating process of $X_{i}$ in our Monte Carlo exercise. We set $n_{t}=100$ for $t=1,\ldots,T$ with $T\in\{4,6,8\}$. We report simulation evidence separately for the two design objectives in (3) and (4), which require different transformation of the outcome variable.
Setup 1: Power maximization. For the power maximization design objective, we compare four design methods: (i) our proposed design; (ii) our design enhanced by the doubly adaptive biased coin design (DBCD) (Hu and Zhang, 2004; Tymofyeyev, Rosenberger and Hu, 2007); (iii) complete randomization; (iv) Neyman allocation. We note that DBCD is a response-adaptive randomization design that more robustly adjusts treatment allocation toward the optimal allocation dynamically. Since our approach involves assigning treatments across multiple covariate strata, we apply the DBCD design separately within each stratum. We also follow the convention in the literature and use the natural logarithm of the viral load as the outcome variable. Following the trial data, the conditional mean and standard deviation of the potential outcomes are $\mu(\mathtt{F},1)=2.50$, $\mu(\mathtt{F},0)=2.98$, $\sigma(\mathtt{F},1)=0.36$, $\sigma(\mathtt{F},0)=2.06$ for the female subgroup, and $\mu(\mathtt{M},1)=2.47$, $\mu(\mathtt{M},0)=2.72$, $\sigma(\mathtt{M},1)=0.82$, $\sigma(\mathtt{M},0)=0.31$ for the male subgroup. Given the calibrated model primitives, the average treatment effect is $\tau=-0.33$. To evaluate the power of different designs under the power maximization objective, we also vary the magnitude of the ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.66$.
@@
Setup 2: Failure reduction. For the failure reduction objective, we again compare across four designs: (i) the proposed method; (ii) our proposed design again enhanced by DBCD; (iii) the ethical design proposed by Rosenberger, Stallard, Ivanova, Harper and Ricks (2001); (iv) complete randomization. In this simulation setup, we transform the original viral load measurements into a binary variable, indicating whether the participant has achieved viral suppression; specifically, if a participant’s viral load is below $1,000$ copies per mL, we set $Y_{i}=1$ (viral suppression). Conditional means of the potential outcomes are $\mu(\mathtt{F},1)=0.78$, $\mu(\mathtt{F},0)=0.57$, $\mu(\mathtt{M},1)=0.84$ and $\mu(\mathtt{M},0)=0.63$. The average treatment effect is $\tau=0.21$. To evaluate the power of different designs under the failure reduction objective, we vary the magnitude of ATE in our data-generating process by perturbing $\mu(\mathtt{M},0)$, such that the perturbed ATE takes values between 0 and $0.42$.
As we discussed in Section 1 and illustrated in Figure 1, the outcomes in the original data exhibit both arm- and covariate-dependent delays, which we also incorporate in our simulation setup. To be precise, the delay variable is generated conditionally according a multinomial distribution: $D_{i}|(X_{i}=x,A_{i}=a)\sim\text{Multinomial}\left(1,\newline \bm{\rho}(x,a)\right)$. Here, we slightly abuse the notation to define $\bm{\rho}(x,a)=(\mathbb{P}(D_{i}=0|x,a),\ldots,\mathbb{P}(D_{i}=T-1|x,a))$, where each element represents the probability of delaying by exactly $d$ stages (instead of cumulative). For $T=4$, the delay mechanism parameters are $\bm{\rho}(\texttt{F},1)=(0.64,0.18,0.07,0.03)$ and $\bm{\rho}(\texttt{F},0)=(0.63,0.18,0.05,0.02)$ for the female subgroup, and $\bm{\rho}(\texttt{M},1)=(0.55,0.23,0.10,0.02)$ and $\bm{\rho}(\texttt{M},0)=(0.54,$ $0.11,0.21,0.01)$ for the male subgroup.
We show in panels (A)–(C) of Figure 3 the optimized variance under the power maximization objective. For complete randomization, the variances of the treatment effect estimator are 7.75, 8.04, and 8.17, respectively. The results clearly demonstrate that the proposed fully forward-looking CARA design and its DBCD-enhanced version exhibit a smaller deviation from the oracle variance (indicated by the horizontal black dashed line); to compare, complete randomization leads to severe efficiency loss. Interestingly, the Neyman allocation performs reasonably well in this design, although it still leads to slightly larger variances.
Figure 3: Optimized variance (panels A–C) and failure rate (panels D–F) for experimental horizons $T=4,5,6$. Covariate stratification is based on biological sex.
Panels (D)–(F) in the figure collects the failure rates for both our proposed method and the ethical design. For complete randomization, the failure rates are 0.31,0.33, and 0.35. Again, our proposed method and its DBCD-enhanced version tend to perform better, and the overall failure rate there is very close to the oracle level (horizontal black dashed line). In contrast, failure rates tend to be considerably higher for complete randomization.
To assess the finite-sample distributional properties of the treatment effect estimator, we present the coverage probabilities in Table 1 for the two design objectives and the three extrapolations methods for the estimated delay mechanism. Encouragingly, the empirical coverage is extremely close to the nominal level. Lastly, we demonstrate and compare the power properties of different design strategies in Figure 4: all methods control the type I error rate well when the true ATE is zero; with $\tau>0$, however, our fully forward-looking CARA design has a clear advantage as it is able to detect the non-zero treatment effect more frequently.
In summary, this synthetic case study showcases that our proposed fully forward-looking CARA design not only provides valid statistical inference in the presence of delayed responses, but it also delivers higher estimation efficiency and enhanced failure reduction for the two design objectives, which is in stark contrast to other methods that overlook the outcome delay issue.
Table 1: Coverage probability comparison.
@@
Conservative 0.94 (0.01) 0.95 (0.01)
Optimistic 0.95 (0.01) 0.94 (0.01)
Neutral 0.95 (0.01) 0.95 (0.01)
@@
Conservative 0.94 (0.01) 0.95 (0.01)
Optimistic 0.96 (0.01) 0.94 (0.01)
Neutral 0.95 (0.01) 0.94 (0.01)
Figure 4: Power comparison for the two design objectives. Covariate stratification is based on biological sex.
Before closing this section, we provide additional simulation evidence with a different covariate stratification, defined by biological sex and the WHO clinical stages of HIV. Specifically, we define four groups: $\texttt{S}_{1}$ for male subjects with WHO stages 1 and 2, $\texttt{S}_{2}$ for female subjects with stages 1 and 2, $\texttt{S}_{3}$ for male with stages 3 and 4, and $\texttt{S}_{4}$ for female with stages 3 and 4.
For the first stratum $\texttt{S}_{1}$, the calibrated parameters are $\sigma(\texttt{S}_{1},1)=1.63$, $\sigma(\texttt{S}_{1},0)=1.87$, $\bm{\rho}(\texttt{S}_{1},1)=(0.10,0.15,0.32,0.04)$, and $\bm{\rho}(\texttt{S}_{1},0)=(0.03,0.10,0.22,0.30)$. For the second stratum $\texttt{S}_{2}$, the parameters are $\sigma(\texttt{S}_{2},1)=1.85$, $\sigma(\texttt{S}_{2},0)=1.90$, $\bm{\rho}(\texttt{S}_{2},1)=(0.14,0.16,0.21,0.05)$ and $\bm{\rho}(\texttt{S}_{2},0)=(0.10,0.14,0.17,0.06)$. For $\texttt{S}_{3}$, we adopt $\sigma(\texttt{S}_{3},1)=2.37$, $\sigma(\texttt{S}_{3},0)=1.63$, $\bm{\rho}(\texttt{S}_{3},1)=(0.10,0.19,0.21,0.14)$ and $\bm{\rho}(\texttt{S}_{3},0)=(0.11,0.16,0.19,0.09)$. And finally for $\texttt{S}_{4}$, we use $\sigma(\texttt{S}_{4},1)=1.73$, $\sigma(\texttt{S}_{4},0)=2.14$, $\bm{\rho}(\texttt{S}_{4},1)=(0.05,0.20,0.17,0.18)$ and $\bm{\rho}(\texttt{S}_{4},0)=(0.19,0.08,0.24,0.02)$. The oracle treatment allocations under the proposed design and the Neyman allocation are demonstrated earlier in Figure 2.
We present simulation results on the optimized variances in panels (A)–(C) of Figure 5, and the failure rates in panels (D)–(F). Although the covariate stratification differs (cf. Figure 3), a consistent pattern emerges: both our proposed design and its DBCD-enhanced variant yield lower variance and enhanced failure reduction by explicitly accounting for the delay mechanism. To compare, the variances of the treatment effect estimator employing complete randomization are 43.96, 45.17, and 47.82, and the failure rates are 0.34, 0.32, and 0.34, respectively. To complement our earlier findings in Figure 4, we also report power comparisons in Figure 6.
Figure 5: Optimized variance (panels A–C) and failure rate (panels D–F) for experimental horizons $T=4,5,6$. Covariate stratification is based on biological sex and WHO clinical stages.
Figure 6: Power comparison for the two design objectives. Covariate stratification is based on biological sex and WHO clinical stages.
6 Conclusion
In this paper, we introduce fully forward-looking covariate-adjusted response-adaptive (CARA) designs that effectively address the challenge of delayed outcomes in adaptive experiments, allowing the delay mechanism to be both arm- and covariate-dependent. Our approach sequentially estimates the delay mechanism, which is then used to inform optimal treatment allocation. Through a comprehensive synthetic case study, we demonstrate that our proposed design can lead to substantial gains in both experimental efficiency and participant welfare compared to traditional CARA designs. We provide rigorous theoretical foundation by deriving semiparametric efficiency bounds in the presence of delayed responses and establishing the consistency and asymptotic normality of treatment effect estimators.